### v1

# VinciCoder: Unifying Multimodal Code Generation via Coarse-to-fine Visual Reinforcement Learning

Abstract
Multimodal code generation has garnered significant interest within the research community. Despite the notable success of recent vision-language models (VLMs) on specialized tasks like Chart-to-code generation, their reliance on single-task training regimens fosters a narrow paradigm that hinders the development of generalized VIsioN Code Intelligence. In this work, we introduce VinciCoder, a unified multimodal code generation model that addresses this limitation via a two-stage training framework. We begin by constructing a large-scale Supervised Finetuning (SFT) corpus comprising 1.6M image-code pairs for tasks involving direct code generation and visual-based code refinement. Subsequently, we introduce a Visual Reinforcement Learning (ViRL) strategy, which employs a coarse-to-fine reward mechanism to improve visual fidelity by calculating visual similarity across local and global image patches. Extensive experiments on various multimodal code generation benchmarks demonstrate that VinciCoder achieves state-of-the-art performance, underscoring the effectiveness of our coarse-to-fine ViRL strategy. The code and model will be available at https://github.com/DocTron-hub/VinciCoder.
footnotetext: Equal contribution.footnotetext: Project leader.footnotetext: Corresponding author.
1 Introduction
Figure 1: VinciCoder is a unified multimodal code generation model built upon the QwenVL series via a two-stage SFT-ViRL training strategy. This approach enables VinciCoder to process visual inputs and generate corresponding code snippets.
Recent advancements in Large Language Models (LLMs) have led to significant breakthroughs in code generation. Leading models, such as Gemini-2.5 [10] and Qwen3-Coder [44], now demonstrate the powerful ability to follow complex textual instructions and generate executable code across multiple languages like Python and Java. Beyond purely textual descriptions, a growing body of research begins to explore multimodal code generation, focusing on visual inputs like charts and web screenshots [38, 54, 17]. These visual modalities are inherently more information-dense than natural language, presenting a more complex challenge for model comprehension and code synthesis.
However, in the multimodal domain, existing works generally focus on training task-specific vision-language models (VLMs). For instance, ChartCoder [54] excels at chart-to-code generation utilizing 160k training samples, while Web2Code [50] constructs a large-scale training corpus for the web-to-HTML domain. Although these specialized models achieve state-of-the-art performance on their respective tasks, their narrow training scopes fundamentally limit their ability to generalize into a unified multimodal code generation framework. Thus, developing a unified multimodal code generation model [16, 35], powered by recent advances in foundation models and reinforcement learning (RL), has emerged as a significant and challenging research direction.
To address these challenges, we introduce VinciCoder, a unified Vision-Language Model (VLM) for multimodal code generation trained via a two-stage SFT-ViRL strategy. For the SFT stage, we construct a large-scale corpus containing 1.3M direct generation samples and 300k examples for a novel task we introduce as code refinement. Distinct from conventional instruction-based editing, this task conditions the model on a target image and a deliberately flawed code snippet that contains logical errors or yields only partial renderings. The model’s objective is to rectify this implementation, generating code whose visual output precisely aligns with the target image. After the SFT phase, we employ RL to further optimize executability and visual fidelity of the generated code. Conventional methods like Reinforcement Learning with Verifiable Rewards (RLVR) are intractable for this domain, as they rely on rule-based textual rewards that are hard to generalize. To overcome this limitation, we propose Visual Reinforcement Learning (ViRL), a framework that pivots the reward mechanism from the textual to the visual domain. ViRL replaces brittle, rule-based validation with a direct reward signal based on perceptual similarity, ensuring scalability and language-agnosticism for the multimodal code generation task.
Furthermore, to robustly compare high-resolution images, our ViRL framework introduces a coarse-to-fine reward function that simultaneously evaluates global structure from downsampled views and local details from segmented image patches. Our experimental results validate this two-stage approach: VinciCoder-SFT, trained on our curated corpus, already establishes a strong baseline, outperforming prior methods on various multimodal code generation benchmarks. The subsequent application of our ViRL strategy then elevates this performance to state-of-the-art (SOTA) levels. To our knowledge, VinciCoder is the first VLM to apply RL for domain-agnostic visual fidelity enhancement of generated code. In summary, our contributions are as follows:
  • We propose VinciCoder, a unified VLM that leverages a two-stage SFT-RL training strategy and visual-based refinement tasks to specialize in diverse multimodal code generation tasks.
  • We propose a novel ViRL framework that employs a coarse-to-fine reward mechanism to precisely evaluate visual similarity between rendered images and their ground-truth targets.
  • We conduct a comprehensive evaluation of VinciCoder across diverse multimodal code generation tasks. The results demonstrate that VinciCoder outperforms existing open-source approaches across benchmarks.
2 Related Works
2.1 MLLMs for Code Generation
Multimodal code generation has received considerable attention, with notable advancements in generating code for visual artefacts such as charts, webpages, scalable vector graphics (SVG), and scientific plots. In the chart domain, previous works [54, 52] focus on constructing large-scale corpora to facilitate SFT. Subsequent works [5, 37] integrate RL to further the fidelity of the generated code. In the web domain, data-centric approaches range from generating synthetic pairs with LLMs [50, 19] to collecting massive real-world datasets [12]. Beyond data construction, recent works like LayoutCoder [42] and LatCoder [13] incorporate layout-aware frameworks to produce more structurally accurate HTML code. Unlike chart and web domains, SVG code offers a rich source of paired data for real-world icons and illustrations. StarVector [28] and OmniSVG [47] leverage this to construct datasets for training on image-to-SVG and text-to-SVG tasks. Similarly, numerous real-world scientific plots and corresponding code are utilized for training. MathCoder-VL [39] leverages the real-world Datikz dataset [4, 3] to train an image-to-code model capable of iteratively converting scientific diagrams into LaTeX code.
However, these methods are typically task-specific, constrained to homogeneous visual patterns and singular code languages. While recent research has shifted towards unified models [16, 35], these efforts rely almost exclusively on SFT, which is insufficient for ensuring either code executability and visual fidelity.
2.2 Multimodal Code Generation Benchmarks
Besides constructing code MLLMs, many multimodal code generation benchmarks have been proposed for evaluation. Previous works generally focus on evaluating direct generation capacity within task-specific domains. For instance, benchmarks such as ChartMimic [45] and Plot2Code [41] evaluate chart-to-code generation capabilities. Other works, including Design2Code [34], UniSVG [21], and Image2Struct [27], assess the code generation of corresponding visual inputs. Also, some benchmarks evaluate beyond direct generation, including interaction [43, 22] and editing generation [7, 53, 46]. Recently, with the growing capacities of VLMs, many new and complex benchmarks have been introduced. Artifactsbench [51] and DCG-Bench [20] propose diverse webpage and chart code generation tasks with dynamic visual images and code complexity.
Figure 2: Our training dataset is constructed via a multi-stage pipeline. We begin by curating a diverse corpus from open-source datasets, employing rigorous filtering and diversity-aware sampling. Subsequently, we enhance the data via two parallel streams: refining existing samples through execution, validation, and optimization, while generating novel ones for the refinement task. This dual strategy yields the final high-quality data pairs for our SFT and RL training.
2.3 RL for MLLM
Inspired by the success of Group Relative Policy Optimization (GRPO) in post-training LLMs [31, 14], RL has garnered significant attention from the research community. Recently, a growing body of work has focused on applying RL to MLLMs to enhance their vision-language reasoning capabilities. Pioneering efforts, such as Vision-R1 [15], VLM-R1 [32], and R1-OneVision [48], first utilize Chain-of-Thought (CoT) data for SFT to establish a reasoning baseline and employ RL to advance the model policy towards generating coherent answers. This RL framework has proven highly versatile, with successful adaptations to a range of fundamental visual tasks, such as grounding [56, 24] and question answering (QA) [36, 6] and segmentation [23]. Besides constructing rewards from generated text contents, another research direction involves leveraging visual feedback to formulate reward functions. For instance, RVRL [8] and RLRF [29] utilize MLLMs and pretrained Vision Transformers (ViTs) to score generated images, thereby providing a reward signal for RL training.
3 Method
3.1 Task Definition
The standard approach for multimodal code generation is to generate code from visual and textual inputs directly. Given an $\operatorname{Image}$ and a $\operatorname{Text}$ instruction, an MLLM is tasked with generating the corresponding $\operatorname{Code}$.
$\operatorname{Code}=\operatorname{MLLM}(\operatorname{Image},\operatorname{Text})$ (1)
Besides direct generation, we introduce a novel visual-based code refinements task, which tasks the model with refining an initial, potentially flawed code draft $\operatorname{Code}_{\operatorname{draft}}$ to the refined version $\operatorname{Code}_{\operatorname{refined}}$.
$\operatorname{Code}_{\operatorname{refined}}=\operatorname{MLLM}(\operatorname{Image},\operatorname{Code}_{\operatorname{draft}},\operatorname{Text})$ (2)
As our experiments demonstrate that the refinement method yields substantial gains over direct generation.
3.2 Data Construction
To construct large-scale multimodal code generation data, we curate datasets from various open-source datasets and generate missing image-code types. All code is re-executed to render the corresponding images. We divide according to image types for illustration.
3.2.1 SFT Data
Chart-to-code. For our chart-to-code task, we curate training data from MCD [16] and MSRL [5]. Our process begins with preprocessing the Chart2Code-160k subset of MCD, where we correct syntax errors using the Ruff linter and custom heuristics. Simultaneously, we curate a diverse 100k-sample subset from the MSRL data using perceptual hashing (pHash) and Mini-Batch K-Means clustering. To construct the code refinement dataset, we first train a specialized chart-to-code VLM on previously curated data. This model is then employed to generate initial code for a new, non-overlapping set of 100k MSRL samples. Acknowledging that the model’s first-pass generations are often imperfect, we form the final refinement dataset by pairing this generated code with the corresponding images and ground-truth targets.
Web-to-HTML. For the web-to-HTML task, we curate training data from MCD [16], Web2M [12], and Web2Code [50]. We first preprocess the Web2M collection, filtering for English-language entries and removing samples with hyperlinks or embedded images. This procedure results in a curated dataset of 60k entries. Following a pipeline similar to our chart-to-code process, we then train a dedicated web-to-code VLM using the combined data from MCD and the filtered Web2M set. Subsequently, this model generates preliminary code for the sampled 100k instances from the Web2Code dataset, forming the final refinement set.
Image-to-SVG. For the image-to-SVG task, our primary training corpus is the ISVGEN subset from UniSVG [21], which contains 360k data pairs. We also construct a refinement dataset with different strategies from the chart and web tasks. Specifically, after training a dedicated VLM on the ISVGEN data, we employ this model to generate code drafts for 100k samples sampled from the ISVGEN training subset itself to constitute the refinement dataset. This strategy is motivated by the observed training instability. The high, fluctuating final training loss suggests the model has not yet memorized the training set, thus validating these samples as non-trivial targets for refinement.
Image-to-Latex. For the image-to-LaTeX task, we curate data from the DaTikZ-v3 [3] and Cosyn-400k [49]. To ensure consistent and valid outputs, we first standardize all code by encapsulating it within a standalone TikZ environment. This step is designed to produce tightly-cropped figures, preventing the generation of full A4-sized PDF pages. Then, we re-execute each sample to validate its integrity, filtering out any instances that result in rendering errors or produce multi-page PDF outputs.
Table 1: Details about scientific plots and corresponding code types in the SFT dataset.
Scientific Plots Code Types Statistics
Document Latex/HTML 71k
Molecule RDKiT/Indigo 50k
Diagram Latex/HTML/Mermaid 48k
Table Latex/HTML 32k
Graphic SVG/Asymptote 27k
Circuit Latex 10k
Scientific Plots-to-code. In addition to the aforementioned domains, we extend our investigation to the challenging area of scientific visualization. This expansion covers a variety of graphical representations, including but not limited to molecular structures, electronic schematics, general diagrams, document layouts, and tabular data. The underlying code for these figures leverages not only the general-purpose languages previously mentioned but also specialized, domain-specific languages such as Mermaid, Asymptote. Our dataset is primarily constructed from the Cosyn-400k collection and various open-source text-to-mermaid datasets. We augment this base with an additional 40k molecular image-code pairs, which we generated by rendering SMILES strings sourced from the USPTO database. Table 1 provides a detailed breakdown of the dataset, summarizing the distribution of image and code types along with their corresponding statistics.
3.2.2 RL Data
For the RL phase, we construct a new dataset spanning five distinct domains, ensuring it is mutually exclusive with our SFT data. The curation process for each domain is as follows: (i) Chart-to-code: We directly utilize the 11k image-code pairs from the second RL-stage subset of the MSRL dataset [5]. (ii) Web-to-HTML: We sample 9k examples from Web2Code [50] and employ Gemini-2.5-Flash [10] to refine the HTML code. This is achieved by prompting the model with varied reference tags and diverse instructions to enhance the visual complexity and quality of the generated webpages. (iii) Image-to-SVG: Using our trained SFT model as a quality filter, we curate the SVG-icons dataset [28], retaining 10k samples that are both high-fidelity and challenging for the model to generate. (iv) Image-to-LaTeX: Leveraging the recently published ImgCode-8.6M dataset [39], we draw samples from its arxiv-woc-680k category. To rectify observed issues such as inconsistent output scaling and low content diversity, we implemented a hybrid filtering pipeline combining both rule-based and model-based approaches, yielding a final, curated set of 10k examples. (v) Chemical Images: We synthesize a new collection of 2k molecular images by rendering unique SMILES strings with both RDKit [18] and Indigo [11].
Table 2: Details about SFT and RL data statistics.
Image Code Data Statistics
SFT Data RL Data
Chart Python 412k 11k
Webpage HTML 355k 9k
Image SVG 463k 10k
Image Latex 108k 10k
Scientific Plots Multi Types 328k 2k
Figure 3: An overview of our coarse-to-fine ViRL strategy. Given an image with instructions, the model generates 8 code rollouts. Each code snippet is first evaluated for a language alignment reward and then rendered into an image. This image is partitioned into local patches (fine-grained) and a downsampled global thumbnail (coarse-grained). The final visual reward is the average cosine similarity between the DINOv2 embeddings of these rendered components and their counterparts from the target image.
The composition of our training datasets is detailed in Table 2. The chart-to-code, web-to-HTML, and image-to-SVG subsets are augmented with 100k, 92k, and 103k refinement data samples, respectively. The remainder of the dataset consists exclusively of direct generation data.
3.3 Model Training
3.3.1 Supervised Finetuning (SFT)
The SFT stage is foundational because our RL phase provides no direct supervisory signal for the textual modality of the code. This initial SFT phase is therefore essential for establishing robust coding capabilities and learning fundamental templates. To this end, we train the model using a standard autoregressive objective.
$\mathcal{L}(\theta):=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{SFT}}}\sum_{t=1}^{T}\log P\left(y_{t}\mid x,y_{<t};\theta\right),$ (3)
$(x,y)$ is the query and target response.
3.3.2 Reinforcement Learning (RL)
Although SFT can yield high performance on token-level metrics, the approach has fundamental limitations for multimodal code generation. First, the autoregressive next-token prediction objective is local in nature. It provides no supervisory signal for global properties such as code executability. Second, the SFT process is not visually grounded, as the model receives no feedback from the rendered output. This is a critical deficiency because the mapping from code to visuals is highly sensitive. Minor modifications in the code can induce substantial changes in the rendered image. This challenge is further compounded across diverse image domains, where factors like resolution can significantly influence perceptual similarity metrics.
To address these limitations, we propose ViRL, a reinforcement learning strategy designed to optimize for visual fidelity. We first render the generated code into an image and then employ the self-supervised ViT, DINOv2 [26], to encode both the rendered and the input source images into visual embeddings. The reward function is subsequently defined as the normalized cosine similarity between these two visual embeddings.
$R_{v}(I_{s},I_{r})=\frac{1}{2}(\cos(\operatorname{DINO}(I_{s}),\operatorname{DINO}(I_{r}))+1)$ (4)
$I_{s}$ and $I_{r}$ are the source and rendered images, respectively.
Coarse-to-fine Visual Reward. Directly downsampling high-resolution images is a flawed approach for computing visual similarity. The process inherently discards crucial fine-grained details, reporting a high similarity by concealing significant visual differences between images. To address this, we propose a coarse-to-fine visual reward mechanism. Given a rendered and a ground-truth image, we first resize the rendered image to match the dimensions of the ground-truth. The fine-grained comparison employs a dynamic tiling strategy that partitions the image into non-overlapping $\times448$ grids. The grid configuration is chosen to match the image’s aspect ratio best. A downsampled thumbnail of the entire image is generated to provide a coarse, global-level context. Subsequently, fine-grained similarity scores are computed between corresponding tiles, while a global similarity score is derived from the thumbnails. The final visual reward is the average of all the scores. This coarse-to-fine mechanism is robust by design. For untiled, low-resolution images, the fine- and coarse-grained rewards converge. The method thus naturally reduces to a standard global comparison, handling all resolutions within a unified framework.
Alignment Reward. We observe that the SFT-trained model sometimes generates code in a language inconsistent with the prompt. To mitigate this, we introduce a code language alignment reward. The reward is computed by parsing the target language, like Python, from the instruction and extracting the language identifier from the generated code, such as ‘‘‘python. We use a predefined mapping to handle aliases, such as mapping tikz to latex. A binary reward of 1 is assigned if the generated language matches a valid alias to enforce instruction fidelity.
Visual Reinforcement Learning (ViRL) During the ViRL stage, we employ the Group Relative Policy Optimization (GRPO) [31] algorithm to fine-tune the model. A primary advantage of GRPO is its independence from a separate critic model, a key component in methods like PPO [30]. For a given input query $x$, the algorithm first samples a set of $G$ responses, $\{o_{1},o_{2},\dots,o_{G}\}$, from the current policy $\pi_{old}$. Each response receives a reward $R_{i}$ and the group-normalized advantage for the $i$-th response at time step $t$ is:
$\hat{A}_{i,t}=\frac{R_{i}-\text{mean}(\{R_{j}\}_{j=1}^{G})}{\text{std}(\{R_{j}\}_{j=1}^{G})}.$ (5)
Table 3: Evaluation results of comparing VinciCoder with various baseline models on multimodal code generation benchmarks. The gray and blue color rows denote the performances of open-source and VinciCoder. The best performance of open-source models is in Bold.
Model ChartMimic_direct_v2 UniSVG-ISVGEN Design2Code Image2Latex_plot ChemDraw
Exec.Rate Low-L High-L Low-L High-L Score Low-L High-L Ren.Succ. EMS Exec.Rate Tani.Sim.
Closed-Source Models
[…]
Open-Source Models
InternVL3-8B 63.3 43.8 46.1 54.5 77.4 68.2 85.3 87.6 57.7 38.6 42.2 6.2
InternVL3.5-8B 66.7 46.0 48.3 55.0 78.0 68.6 85.8 87.3 58.3 40.5 49.2 7.8
InternVL3-14B 72.3 51.3 54.1 51.4 75.5 65.8 85.8 87.5 73.3 52.2 71.1 40.2
InternVL3.5-14B 73.2 52.8 55.4 52.0 75.0 65.9 86.1 87.8 73.0 50.2 71.9 39.3
Qwen3-VL-32B 83.0 66.9 77.5 68.0 86.0 78.8 88.6 89.8 75.7 53.3 37.5 48.8
InternVL3.5-38B 79.0 60.0 71.8 51.9 77.3 67.1 87.8 88.4 72.6 49.5 55.5 31.4
Qwen2.5-VL-72B 88.5 72.7 79.1 47.7 76.0 64.7 86.9 88.7 62.0 41.7 75.8 28.0
Qwen2.5-VL-7B 68.7 42.2 40.1 47.5 73.8 63.3 83.4 87.6 42.7 25.5 21.1 11.7
[…]
Qwen3-VL-8B 78.3 62.5 67.8 53.0 77.0 67.4 85.5 87.2 47.7 33.0 78.9 41.2
[…]
It then optimizes the new policy $\pi_{\theta}$ by maximizing the following objective function:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D},\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x)\end{subarray}}$ (6)
$\biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\biggl(\min\Bigl(r_{i,t}(\theta)\hat{A}_{i,t}$
$\operatorname{clip}\left(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Bigl)\biggr]$
where the probability ratio $r_{i,t}(\theta)$ is defined as:
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$ (7)
The reward $R$ is the combination of coarse-to-fine visual reward $R_{v}$ and language alignment reward $R_{l}$, which is calculated as
$R_{i}=\omega_{v}\frac{1}{N}\sum_{i=1}^{N}R_{i,v}+\omega_{l}R_{l}.$ (8)
When the generated code fails to render images, the $R_{v}$ will be set to 0. The $\omega_{v}$ and $\omega_{l}$ are set to 0.9 and 0.1, respectively. The framework of our proposed coarse-to-fine reward mechanism and ViRL strategy are denoted in Figure 3.
4 Experiments
4.1 Implementation Details
We conduct SFT on this dataset for one epoch, using Qwen2.5-VL-7B-Instruct [2] and Qwen3-VL-8B-Instruct as base models. The training is performed on 24 H800 GPUs with a global batch size of 96. In the RL stage, we utilize GRPO to train the SFT-model , allocating 16 GPUs for the policy model and 4 GPUs for reward scoring, with a global batch size of 256. The DINOv2-L is utilized for visual embedding. The SFT and RL training phases are implemented using the ms-swift [55]and Easyr1 [33] frameworks.
4.2 Evaluation Settings
We conduct comprehensive experiments to evaluate the unified multimodal code generation capabilities of VinciCoder and baseline models, including closed-source [10, 1, 25] and open-source models [2, 57, 40]. Our evaluation spans five domains, four of which utilize established benchmarks: ChartMimic [45] direct generation for chart-to-code, Design2Code [34] for webpage-to-HTML, UniSVG [21] ISVGEN set for image-to-SVG, and Image2Struct [27] for image-to-LaTeX generation. For the UniSVG, the final score is a weighted combination of two components. The low-level score, which is the average of SSIM and (1 - LPIPS), contributes 40% to the final score, while the high-level score contributes the remaining 60%. For the Image2Struct, we utilize Earth Mover Similarity (EMS) in their GitHub repository for evaluation. For the molecule-to-code task, we evaluate performance on the Cosyn-400k [49] chemistry test set, which consists of 128 molecular images paired with their ground-truth SMILES strings. We report two metrics: the execution rate and the average Tanimoto similarity (Tani. Sim.) of the generated SMILES.
Figure 4: The reward progression during our ViRL training stage. The learning curves illustrate that as training progresses, the visual reward steadily increases, while the alignment reward rapidly converges to and then plateaus at its maximum value of 1.
4.3 Main Results
As detailed in Table 3, our comparative analysis concludes that VinciCoder sets a new state-of-the-art among open-source solutions. It outperforms existing models across the vast majority of benchmarks, with a particularly significant lead over competitors of a similar scale. The comparison against proprietary models further demonstrates the superior performance of our proposed VinciCoder on tasks like image-to-SVG and ChemDraw, yet leading closed-source commercial models maintain an advantage in areas such as chart-to-code and webpage-to-HTML.
The comparison between SFT and the final models further validate the effectiveness of our two-stage training process. First, the SFT stage alone yields substantial gains, with VinciCoder-SFT significantly outperforming its base model across nearly all benchmarks. This result underscores the high quality and efficacy of our large-scale SFT dataset. The subsequent ViRL stage further enhances performance across almost all metrics, particularly visual similarity and execution rate. This improvement stems from two distinct mechanisms. The execution rate increases because the RL penalizes invalid outputs by assigning a zero reward to non-executable code. Concurrently, the ViRL strategy drives the enhanced visual similarity by directly optimizing the perceptual alignment between rendered and target images. We visualize the reward progressions during the ViRL training stage in Figure 4.
In addition to generalist models, we conduct a comparative analysis against task-specific baselines in the Appendix.
4.4 Ablation Studies
Training Strategy and Data Scales. To validate our SFT-RL training procedure, we conduct an ablation study on the Qwen2.5-VL model, analyzing the impact of each stage by varying its respective data scales.
Figure 5: The ablation study about SFT and RL training stage.
We first ablate the contributions of the SFT and RL stages. As illustrated in Figure 5, the two-stage SFT-RL strategy consistently yields the best performance. This result aligns with the established paradigm where SFT provides a strong initial policy that is subsequently refined via targeted RL. Furthermore, applying our ViRL strategy directly to the base model (without SFT) also yields demonstrable improvements. On the ChartMimic benchmark, ViRL leads to significant gains in execution rate and high-level scores, indicating that its visual feedback effectively enhances code correctness and perceptual fidelity. Also, the improvement of the low-level score further demonstrates that ViRL optimize the generated code for the visual feedback. However, the performance gains on the UniSVG are marginal. This suggests that while ViRL is effective, its efficacy is inevitably contingent upon the foundational code generation capacity of the base model.
Figure 6: The ablation experiments about model performance under various SFT data scales.
Figure 7: Showcases of images rendered by VinciCoder-generated code and other similar scales open-source VLMs.
Furthermore, we investigate the data scaling during the SFT stage. As illustrated in Figure 6, the results indicate that model performance scales predictably with the size of the SFT dataset, showing an initial upward trend that gradually saturates. This observation validates the utility of our SFT data but also demonstrates that performance saturates, underscoring the necessity of RL for further improvement.
Table 4: Ablation studies about utilizing various metrics as the reward function and our proposed coarse-to-fine reward. InternViT is the abbreviation of InternViT-300M-448px.
Ablations ChartMimic UniSVG Design2Code
Low-L High-L Low-L High-L Low-L High-L
Reward Formulations
SSIM 75.7 75.3 77.4 89.8 84.5 85.6
InternViT 77.9 78.1 77.1 90.6 86.3 87.9
Coarse-to-fine ViRL
w/o Coarse 77.8 78.9 76.9 91.8 87.4 88.3
w/o Fine 77.0 78.0 77.1 91.8 86.9 87.6
w/o Align 78.2 79.4 76.2 90.5 87.5 88.3
DINO-L 78.3 79.8 77.0 92.0 88.2 89.1
Reward Functions. We conduct several ablation studies to justify the design of our visual reward, which combines a coarse-to-fine strategy with DINOv2-L embeddings. All studies use VinciCoder-SFT-7B as the base model. First, we investigate alternative reward formulations, comparing our approach against a low-level perceptual metric (SSIM) and a semantic-level reward from the image-text aligned ViT (InternViT-300M-448px-V2.5 [9]). We select InternViT as its parameter count is the same as DINOv2-L, which allows for a controlled comparison. In addition, we further ablate the individual components of our coarse-to-fine reward strategy to isolate their impact. The results in Table 4 confirm that our proposed coarse-to-fine reward achieves the most balanced and effective performance. Our analysis indicates that self-supervised DINOv2 excel at capturing the fine-grained features essential for visual similarity evaluation, in contrast to image-text aligned ViTs, which are optimized for global semantic alignment. The ablation study also highlights the critical role of our coarse-to-fine strategy for high-resolution images. The performance degrades significantly without both reward components on the ChartMimic and Design2Code benchmarks. Conversely, performance on the lower-resolution UniSVG benchmark is stable, as its images are small enough that our method naturally reduces to a single global comparison, making the coarse and fine-grained rewards equivalent.
Refinement Tasks. Our experiments reveal that incorporating the refinement data into the SFT stage measurably boosts model performance. We conduct an ablation study to evaluate the impact of this data by directly comparing the performance of models trained with and without it. The results in Table 5 show that incorporating the refinement data leads to improvements across nearly all metrics, demonstrating the effectiveness of this data.
Table 5: Ablation studies about the 300k refinement data. We compare models after the SFT stage.
Model ChartMimic UniSVG Design2Code
Low-L High-L Low-L High-L Low-L High-L
w/o Refine 76.2 76.4 75.6 89.4 86.1 86.6
w/ Refine 75.8 78.6 78.2 91.0 86.5 87.0
5 Conclusion
We introduce VinciCoder, a unified multimodal code generation model trained via a two-stage SFT-ViRL strategy in this work. We construct a large-scale dataset of 1.6M samples from various domains to perform SFT. For the subsequent RL stage, we propose a novel coarse-to-fine ViRL training strategy, which calculates visual similarity rewards from both coarse-grained and fine-grained image patches. Experimental results demonstrate the superior performance of VinciCoder, validating the effectiveness of our proposed unified multimodal code generation model and the ViRL training paradigm.
Supplementary Material
5.1 Benchmark Details
Our evaluation methodology for the ChartMimic, UniSVG, and Design2Code benchmarks follows the official implementations provided in their respective GitHub repositories. Regarding Image2Struct, we observed a minor inconsistency between the metrics in the published paper and the official code. To ensure replicability, we follow the implementation in the GitHub repository. We introduce the our calculation of EMS in Section 5.1.1. For the ChemDraw benchmark, we evaluate performance using the execution rate and Tanimoto similarity, two metrics we introduce in Section 5.2.
5.1.1 Earth Mover’s Similarity (EMS)
The Earth Mover’s Similarity (EMS) is an efficient, patch-based metric derived from the Earth Mover’s Distance (EMD). It quantifies similarity through a two-level process that compares the arrangement of image patches at a global level and the pixel distributions within those patches at a local level.
The calculation begins by converting the reference image $x$ and the generated image $\hat{x}$ to grayscale. The most frequent pixel value in the reference image is identified and designated as the background color, $v_{bg}$, to focus the analysis on foreground content.
Next, both images are partitioned into a grid of $K$ patches, $\{P_{0},\dots,P_{K-1}\}$. A global signature, $S_{global}=\{(w_{i},P_{i},c_{i})\}_{i=0}^{K-1}$, is created for each image. Each element in the signature represents a patch $P_{i}$ by its content, its normalized center coordinate $c_{i}$, and a weight $w_{i}$. To prioritize meaningful content, this weight $w_{i}$ is set significantly higher for patches containing non-background pixels.
The dissimilarity between the images is calculated using a hierarchical cost matrix $C_{p}$, where each element $C_{p}[i,j]$ defines the cost of matching patch $P_{i}$ from image $x$ to patch $P_{j}$ from image $\hat{x}$. This cost aggregates two factors: the dissimilarity within the patches and the spatial distance between them:
[…]
Algorithm 1 Tanimoto Similarity Calculation
1:function CalculateTanimoto(predicted_smiles, ground_truth_smiles)
[…]
3:   $\textit{pred\_mol}\leftarrow\text{RDKit.MolFromSmiles}(\textit{predicted\_smiles})$
4:   $\textit{gt\_mol}\leftarrow\text{RDKit.MolFromSmiles}(\textit{ground\_truth\_smiles})$ $\triangleright$ Step 2: Handle invalid SMILES. If either is invalid, similarity is 0.
5:   if pred_mol is null or gt_mol is null then
6:     return $0.0$
7:   end if
[…]
9:   $\textit{pred\_fp}\leftarrow\text{GetMorganFP}(\textit{pred\_mol},\text{radius}=2)$
10:   $\textit{gt\_fp}\leftarrow\text{GetMorganFP}(\textit{gt\_mol},\text{radius}=2)$
[…]
12:   $\textit{similarity}\leftarrow\text{TanimotoSimilarity}(\textit{pred\_fp},\textit{gt\_fp})$
13:   return similarity
14:end function
Here, $\text{EMD}_{intra}$ is the classic EMD computed on the pixel values and their local coordinates within each patch. The term $c_{i}-c_{j}||_{1}$ is the $L_{1}$ (Manhattan) distance between the patch centers, scaled by a factor $\lambda$:
$\lambda=\frac{\sqrt{r\times s}}{W+H}$ (10)
where $(r,s)$ are the patch dimensions and $(W,H)$ are the image dimensions.
[…]
$\text{EMS}(x,\hat{x})=\max\left(0,1-\frac{\text{EMD}_{block}(x,\hat{x})}{\text{EMD}_{block}(x,x_{const})}\right)$ (11)
An EMS score of 1 indicates identical images, while a score of 0 indicates maximum dissimilarity.
5.2 ChemDraw Metrics
Performance on the molecule-to-code task is assessed using two key metrics. The first, execution rate, serves as a primary filter for syntactic correctness, evaluating whether the VLM produces chemically valid code and SMILES strings. The second, Tanimoto similarity, measures the chemical fidelity of the valid generations. This metric utilizes the RDKit library to compare the molecular fingerprints of the generated structure against the ground-truth structure. A detailed description of the calculation in Algorithm 1.
The formula for Tanimoto similarity between two fingerprints (bit vectors) $A$ and $B$ is:
$T(A,B)=\frac{|A\cap B|}{|A\cup B|}=\frac{c}{a+b-c}$
where:
  • $a$ is the number of bits set in fingerprint $A$.
  • $b$ is the number of bits set in fingerprint $B$.
  • $c$ is the number of bits set in both $A$ and $B$ (the intersection).
6 More Experiment Results
6.1 Comparison with task-specific models
We compare VinciCoder with other domain-specific models, like chart-to-code and image-to-SVG models. The results in Table 6 show that VinciCoder performs comparably to specialized models such as MSRL [5] and ChartMaster [37] on the ChartMimic dataset. More domain-specific models are compared in the Appendix. Moreover, its performance on UniSVG exceeds that of fine-tuned specialist models, which validates the effectiveness of our unified framework for multimodal code generation.
Table 6: Comparing with task-specific chart-to-code models.
Chart-to-code Models ChartMimic_direct_v2
Exec. Rate Low-Level High-Level
ChartCoder [54] 91.4 72.5 74.0
Chart2Code[52] 62.1 42.9 33.3
MSRL [5] 96.5 78.6 83.8
ChartMaster [37] 93.8 78.2 85.1
VinciCoder-7B 91.2 78.3 79.8
VinciCoder-8B 91.6 78.9 80.6
Table 7: Comparing with task-specific image-to-SVG models. The metrics of finetuned models are from UniSVG.
Image-to-SVG Models UniSVG-ISVGEN
SSIM $\uparrow$ LPIPS $\downarrow$ CLIP $\uparrow Score$
LLaVA 1.5 Tuned 65.4 47.9 80.2 71.6
Llama 3.2 Tuned 72.2 37.8 84.3 77.5
Qwen2.5-VL Tuned 72.5 36.8 83.6 77.3
VinciCoder-7B 77.4 23.3 92.0 86.0
VinciCoder-8B 77.4 23.2 94.1 87.3
6.2 ViRL Results
Figure 4 shows the improvement in visual similarity during RL training. Our reward function assigns a score based on visual similarity only for executable rollouts, while all unexecutable rollouts receive a reward of zero. We want to analyze whether the visual similarity between executable rollouts and input images improves during RL training. We visualize the visual reward of executable rollouts only during the RL training process. The result in Figure 8 denotes that during our training, the visual similarity of executable rollouts improves as well, demonstrating that our proposed ViRL training strategy not only improves the execution rate but also the visual fidelity.
7 Evaluation Prompts
In our experiments, all the prompts utilize the official implementations, except for our constructed ChemDraw. We visualize the prompt and the input molecule image in Figure 9.
Figure 8: The reward progression of executable rollouts during our ViRL training stage.
ChemDraw Benchmark Prompt Instruction: Create a python script that displays a visual representation of the molecule shown in the image. Utilize ’smiles = <<SMILES>> string’ to define the smile and RKiT or Indigo for plot. Image:Refer to caption
Figure 9: Prompt for our constructed ChemDraw benchmark.